OpenStudy (eric_d):

Factorise b) 2(8 factorial)-7 factorial @mathmale

OpenStudy (mathmale):

Is that equivalent to 2*8!-7! ?

OpenStudy (eric_d):

yes

OpenStudy (mathmale):

Please factor 8! by factoring out the 8. What's left?

OpenStudy (eric_d):

1

OpenStudy (mathmale):

Hint: 8! = 8*7*6*5*4*3*2*1 What would happen if you were to divide both sides by 8?

OpenStudy (eric_d):

The whole equation

OpenStudy (eric_d):

@mathmale

OpenStudy (unklerhaukus):

(n+1)! = (n+1) n!

OpenStudy (mathmale):

Eric, I'm sorry, but OpenStudy was not responding for a while. Your job here is to find the factors of 2*8! - 7! I've been asking you to focus on that 8! alone. 8! = 8*7!, so 2*8! - 7! = 2*8*7! - 7! Can you factor that?

OpenStudy (eric_d):

I got it... It's 15(7!)

OpenStudy (eric_d):

Thanks @mathmale

OpenStudy (mathmale):

My great pleasure. Always happy to be in touch with you.